OpenStudy (ovoxoclovd):

A scientist looks at a bacterium and a virus in a lab. The bacterium has a diameter of 10^-6 meters. The virus has a diameter of 10^-7 meters. Which statement accurately compares the sizes of the specimens? The diameter of the bacterium is 10 times greater than that of the virus. The diameter of the bacterium is 1/100 times as great as that of the virus. The diameter of the virus is 10 times greater than that of the bacterium. The diameter of the virus is 1/100 times as great as that of the bacterium.

OpenStudy (thecatman):

take a guess so i have something teach from

OpenStudy (thecatman):

only hint i have is that its not c

OpenStudy (ovoxoclovd):

I thought it was c lol

OpenStudy (thecatman):

pay attention to the 10^-6 and the 10^-7

OpenStudy (thecatman):

10^-6 is larger than 10^-7 and its up to you to find out how mutch

OpenStudy (thecatman):

.000001 vs .0000007

OpenStudy (thecatman):

.0000001 sorry not .0000007

OpenStudy (thecatman):

whats the difference between bacterium (.000001) and virus (.0000001) ?

OpenStudy (ovoxoclovd):

There is no difference ugh i hate math

OpenStudy (thecatman):

its not 1/100 of any and its not c ;)

OpenStudy (ovoxoclovd):

Is it A?

OpenStudy (thecatman):

correct

OpenStudy (ovoxoclovd):

YAY OMG THANK YOU

OpenStudy (thecatman):

np

OpenStudy (thecatman):

thank you

jimthompson5910 (jim_thompson5910):

You can divide the two values A = diameter of bacterium B = diameter of virus A/B = (10^(-6))/(10^(-7)) A/B = 10^(-6-(-7)) A/B = 10^(-6+7) A/B = 10^(1) A/B = 10 Since the ratio of the two diameters is 10, this means that A is ten times bigger

OpenStudy (ovoxoclovd):

Thank you (:

OpenStudy (magallar23):

Answer is A